PROSPECTS FOR PRECISION MEASUREMENTS ON AMMONIA MOLECULES IN A FOUNTAIN

\underline{HENDRICK L. BETHLEM}, Laser Centre Vrije Universiteit, De Boelelaan 1081, NL-1081HV Amsterdam, The Netherlands.

The recent demonstration of cooling and manipulation techniques for molecules offers new possibilities for precision measurements in molecules. At the LCVU, we are constructing a molecular fountain based on a Stark decelerated molecular beam. In this fountain, ammonia molecules are decelerated, cooled, and subsequently launched upwards some 10-50 cm before falling back under gravity, thereby passing a microwave cavity or laser beam twice – as they fly up and as they fall back down. In this way, it will be possible to measure the inversion transition in ^{15}NH_{3} around 22.6 GHz with an fractional accuracy better than 10^{−12} ^{a}. Besides serving as a proof-of-principle, these measurements may be used as a test of the time-variation of fundamental constants using the sensitivity of the tunneling motion to a change of the proton-electron mass ratio. We present first experimental results of the molecular beam decelerator and focusing lenses.

aHendrick L. Bethlem, Masatoshi Kajita, Boris Sartakov, Gerard Meijer, and Wim Ubachs, to appear in Eur. Phys. J. Special Topics.